[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jun 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 26 02:32:29 UTC 2017


HSU Shift Log for Sun Jun 25 2017 22:32:21 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: GOES-16 imagery stopped in NAWIPS around 20z.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170626/1b8ba6b1/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list